z^2/2 = z/5 + 7/10
Answer by mukhopadhyay(490)

z^2/2 = z/5 + 7/10
=> z^2/2 = (2z+7)/10
=> 5z^2 = 2z+7
=> 5z^2 - 2z - 7 = 0
=> 5z^2 - 7z + 5z - 7 = 0
=> (5z+1)(z-7) = 0
=> z=-1/5 or z=7